OpenStudy (anonymous):

Fully factorize: x^2 - a^2 +x +a Can anyone solve this for me step by step?

OpenStudy (zzr0ck3r):

\(x^2-a^2+x+a=(x-a)(x+a)+(x+a)=(x+a)(x-a+1)\)

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

let me get off my tablet, brb

OpenStudy (anonymous):

Okay

OpenStudy (freckles):

So you are to factor x^2-a^2+x+a?

OpenStudy (anonymous):

@freckles yupp

OpenStudy (freckles):

\[x^2-a^2=(x-a)(x+a) \\ \text{ So} (x-a)(x+a)+1(x+a)=(x+a)(x-a+1)\] Like Zocker said.. \[(x-1+a)(x-1-a)=(x-1)^2-a^2=x^2-2x+1-a^2 \\=x^2-a^2-2x+1 \neq x^2-a^2+x+a\]

OpenStudy (anonymous):

Yeah, you both are right. Turns out, I was looking at the wrong answer in my textbook and that's the right one, Thanks, though. :)